# Learning to Adapt SAM for Segmenting Cross-domain Point Clouds

### v3

@@ Abstract
Unsupervised domain adaptation (UDA) in 3D segmentation tasks presents a formidable challenge, primarily stemming from the sparse and unordered nature of point cloud data. Especially for LiDAR point clouds, the domain discrepancy becomes obvious across varying capture scenes, fluctuating weather conditions, and the diverse array of LiDAR devices in use. While previous UDA methodologies have often sought to mitigate this gap by aligning features between source and target domains, this approach falls short when applied to 3D segmentation due to the substantial domain variations. Inspired by the remarkable generalization capabilities exhibited by the vision foundation model, SAM, in the realm of image segmentation, our approach leverages the wealth of general knowledge embedded within SAM to unify feature representations across diverse 3D domains and further solves the 3D domain adaptation problem. Specifically, we harness the corresponding images associated with point clouds to facilitate knowledge transfer and propose an innovative hybrid feature augmentation methodology, which significantly enhances the alignment between the 3D feature space and SAM’s feature space, operating at both the scene and instance levels. Our method is evaluated on many widely-recognized datasets and achieves state-of-the-art performance.
@@ 1 Introduction
Figure 1: Comparison of 3D UDA paradigms. Different from aligning two feature domains directly, our method makes both the source domain and target domain align with the SAM feature space.
@@
Considering that RGB cameras yield dense, color-rich, and structured data, and more importantly, they represent minor discrepancies across various devices, certain 3D UDA methods [19, 5, 4] utilize the synergy of LiDAR and camera capabilities to achieve more comprehensive and precise perception, and further enhance adaptation capabilities for 3D segmentation tasks. However, these methods usually train 2D and 3D networks simultaneously demanding substantial online computing resources.
Vision foundation models (VFMs), such as the Segment Anything Model (SAM) [23], have garnered significant attention due to their remarkable performance in addressing open-world vision tasks. Such models are trained on massive image data with tremendous parameters. Compared with a common model trained on limited data, VFMs have more general knowledge and much stronger generalization capability. Many works such as  [8, 7] have emerged recently to transfer the general 2D vision knowledge of VFMs to 3D and have achieved promising performance.
@@
Based on SAM, focusing on image segmentation, we propose a novel paradigm for 3D UDA segmentation. As shown in Fig. 1, different from previous UDA approaches that strive to align the target domain to the source domain so that the model trained on labeled source data can also work on target data without annotation, our method makes both the source domain and target domain align with the SAM feature space. SAM feature space contains more general knowledge, which provides a friendly space to unify the feature representation from different domains. We utilize RGB images to assist point clouds in our framework. However, unlike methods mentioned above only using images to provide auxiliary information, we take images as a bridge to align diverse 3D feature spaces to the SAM feature space, so we do not need to train extra 2D networks and can process the image offline for less computing resources. Considering that the 3D feature space created by the source-domain data and the target-domain data is still much smaller than the SAM feature space, we propose a hybrid feature augmentation method at both scene and instance levels to generate more 3D data with diverse feature patterns in a broader data domain, which can further benefit the 3D-to-SAM feature alignment. In particular, we make full use of the masks generated by SAM to mix instance-level point clouds with the other domains. This technique can maintain the geometric completeness of instances, which is beneficial for semantic recognition.
@@
To verify the effectiveness of our method, we compare it with current SOTA works on extensive 3D UDA segmentation settings and our method outperforms others by a large margin, improving about $14\%$ mIoU for VirtualKITTI-to-SematicKITTI, about $15\%$ mIoU for Waymo-to-nuScenes, and about $20\%$ mIoU for nuScenes-to-SemanticKITTI domain adaptation. Surprisingly, our unsupervised method achieves comparable performance with the supervised method for city-changing and light-changing settings on the nuScenes dataset. Furthermore, we also test our method on more challenging tasks, such as panoptic segmentation and domain generalization, and experimental results show that our method is robust and has good generalization capability.
@@ 2.2 Domain Adaptation for 3D Segmentation
Unsupervised Domain Adaptation (UDA) aims at transferring knowledge learned from a source annotated domain to a target unlabelled domain, and there are already several UDA methods proposed for 2D segmentation [6, 46, 22, 49]. In recent years, domain adaptation techniques have gained increasing traction in the context of 3D segmentation tasks. [45] leverage a ”Complete and Label” strategy to enhance semantic segmentation of LiDAR point clouds by recovering underlying surfaces and facilitating the transfer of semantic labels across varying LiDAR sensor domains. CosMix [36] introduces a sample mixing approach for UDA in 3D segmentation, which stands as the pioneering UDA approach utilizing sample mixing to alleviate domain shift. It generates two new intermediate domains of composite point clouds through a novel mixing strategy applied at the input level, mitigating domain discrepancies. However, due to the sparsity and irregularity of the point cloud, the disparity across different point cloud data domains is larger compared to that across 2D image domains, which makes it difficult to mitigate the variation across domains.
@@ 2.3 Vision Foundation Models
The rise of foundation models [12, 21, 39] has garnered significant attention which are trained on extensive datasets, consequently demonstrating exceptional performance. Foundation models [48, 41] have seen significant advancements in the realm of 2D vision, and several research studies extend these foundation models to comprehend 3D information. Representative works CLIP [35] leverage contrastive learning techniques to train both text and image encoders, CLIP2Scene [8] extends the capabilities of CLIP by incorporating a 2D-3D calibration matrix, facilitating a deeper comprehension of 3D scenes. The Meta Research team recently launched the ‘Segment Anything Model’ [23], trained on an extensive dataset of over 1 billion masks from 11 million images. Utilizing efficient prompting, SAM can generate high-quality masks for image instance segmentation. The integration of flexible prompting and ambiguity awareness enables SAM with robust generalization capabilities for various downstream segmentation challenges. Many methods [7, 8, 28] take it as an off-the-shelf tool and distillate the knowledge to solve 3D problems by 2D-3D feature alignment. In our work of tackling the UDA of 3D segmentation, we utilize SAM to provide 2D prior knowledge for 3D feature alignment in a wider data domain.
@@ 3 Method
Figure 2: Pipeline of our method. The point cloud is fed into the point encoder for point embeddings at the top, and the corresponding images are passed through the SAM encoder for image embeddings at the bottom, from which we obtain SAM-guided point embedding with the 2D-3D projection. Alignment loss $L_{align}$ is calculated based on the SAM-guided features and original features. Furthermore, augmented inputs by mixing source and target data, providing diverse feature patterns, are used to boost the 3D-to-SAM feature alignment.
@@ 3.1 Problem Statement
We explore UDA for 3D segmentation, in which we have the source domain, denoted as $D_{S}=\{P_{S},I_{S},Y_{S}\}$ with paired input, namely point cloud $P_{S}$ and image $I_{S}$, as well as annotated labels $Y_{S}$ for each point, and the target domain denoted as $D_{T}=\{P_{T},I_{T}\}$ without any annotation. Using these data, we train a 3D segmentation model that can generalize well to the target domain.
3D data from various domains have obvious differences in distribution and patterns, leading to over-fitting problems when models trained in one domain try to analyze data from another. The main solution is to align different features despite domain differences to achieve the generalization capability of the model. Different from previous works, our novel paradigm is to map data from distinct domains into a unified feature space, ensuring the model performs consistently across domains.
@@ 3.2 Framework Overview
The vision foundation model, SAM, is trained by massive image data, which contains general vision knowledge and provides a friendly feature space to unify diverse feature representations. Taking 2D images as the bridge, the 3D feature space of different domains can be indirectly unified by bringing them closer to the SAM feature space based on 2D-to-3D knowledge distillation. Based on this, we design a novel SAM-guided UDA method for 3D segmentation, as Fig. 2 shows.
Specifically, given a point cloud input $P$, the point encoder $M$ generates a point embedding $F_{point}\in\mathbb{R}^{n\times d}$ in the $d$-dimensional latent feature space. Concurrently, the corresponding image input $I$ is passed through the SAM encoder for a c-channels image embedding $F_{image}\in\mathbb{R}^{h\times w\times c}$. Utilizing the correspondence between the point cloud and image, we acquire SAM-guided point embedding $\hat{F}_{point}\in\mathbb{R}^{n\times d}$ to compute the alignment loss $L_{align}$ with the original point embedding $F_{point}$, serving the purpose of using SAM as a bridge to integrate the features of diverse data domains into a unified feature space. Notably, during training, the input for feature alignment consists of data from both source and target domains. We named this process as SAM-guided Feature Alignment. At the same time, as for labeled data $Y$, segmentation loss $L_{seg}$ is also calculated as semantic supervision. During model training, only the point cloud branch of the whole pipeline is trained, and the gradient is not calculated in the image branch, which makes our method more lightweight. Furthermore, a Scene-Instance Hybrid Feature Augmentation is designed, which consists of scene-level and instance-level mix-up strategy. These mix-up strategies boost the variance of training data and generalize the network capability under the convex combination of the source domain and target domain data. Notably, the instance-level feature augmentation could maintain the local geometric relationship between two domains and make the subsequent alignment efficient.
@@ 4.1 Dataset Setup
We first follow the benchmark introduced in xMUDA [20] to evaluate our method, comprehending four domain shift scenarios, including (1) USA-to-Singapore, (2) Day-to-Night, (3) VirtualKITTI-to-SemanticKITTI and (4) A2D2-to-SemanticKITTI. The first two leverage nuScenes [3] as their dataset, consisting of 1000 driving scenes in total with 40k annotated point-wise frames. Specifically, The former differs in the layout and infrastructure while the latter exhibits severe illumination changes between the source and the target domain. The third is more challenging since it is the adaptation from synthetic to real data, implemented by adapting from VirtualKITTI [14] to SemanticKITTI [2] while the fourth involves A2D2 [15] and SemanticKITTI as different data domains, where the domain discrepancy lies in the distinct density and arrangement of 3D point clouds captured by different devices since the A2D2 is captured by 16-beam LiDAR and the SemanticKITTI uses 64-beam LiDAR. For the above settings, noted that only 3D points visible from the camera are used for training and testing, specifically, only one image and corresponding points for each sample are used for training. Since we only use the image combined with SAM as offline assistance for the training of a 3D segmentation network instead of training a new 2D segmentation network, we focus on comparing the performance of the 3D segmentation network and enable model training with the whole point cloud sample because of less computational cost, even if some part of it is not visible in the images. For the part of the point cloud that cannot be covered by the image, alignment loss is not calculated, only segmentation loss is calculated. Thus, we also compare our method with others trained with the whole ${}^{\circ}$ view of the point cloud, in which three datasets are involved including nuScenes, SemanticKITTI, and Waymo [29]. In these settings, we use 6 images in nuScenes covering ${}^{\circ}$ view, 1 image in SemanticKITTI covering ${}^{\circ}$ view, and 5 images in Waymo covering ${}^{\circ}$ view. More information is introduced in the Appendix. For metric, We compute the Intersection over the Union per class and report the mean Intersection over the Union (mIoU).
@@ 4.2 Implementation Details
We make source and target labels compatible across these experiments. For all benchmarks in prior multi-modal UDA methods, we strictly follow class mapping like xMUDA for a fair comparison, while we map the labels of the dataset in other experiments into 10 segmentation classes in common. Our method is implemented by using the public PyTorch [31] repository MMDetection3D [11] and all the models are trained on a single 24GB GeForce RTX 3090 GPU. To compare fairly, we use SparseConvNet [16] with U-Net architecture as the 3D backbone network when following the benchmark introduced in xMUDA to evaluate our method and use MinkUNet32 [9] as the 3D backbone network when following the setup of taking the whole ${}^{\circ}$ point cloud as input, which is also the backbone of the state-of-the-art uni-modal method CosMix. For the image branch, the ViT-h variant SAM model is utilized to generate image embedding for SAM-guided feature alignment and instance masks for hybrid feature augmentation in an offline manner. We keep the proportion of mixed data and normal data from the source and target domain the same during model training. Before the data is fed into the 3D network, data augmentation such as vertical axis flipping, random scaling, and random 3D rotations are widely used like all the compared methods. For the model training strategies, we choose a batch size of 8 for both source data and target data, then mix the data batch for training at each iteration. Besides, we adopt AdamW as the model optimizer and One Cycle Policy as the learning-rate scheduler.
@@
Method USA normal-$\rightarrow$ Singapore Day normal-$\rightarrow$ Night v.KITTI normal-$\rightarrow$ Sem.KITTI A2D2 normal-$\rightarrow$ Sem.KITTI
Source only
$62.8$
$+0.0$
$68.8$
$+0.0$
$42.0$
$+0.0$
$35.9$
$+0.0$
xMUDA [20]
$63.2$
$+0.4$
$69.2$
$+0.4$
$46.7$
$+4.7$
$46.0$
$+10.1$
DsCML [32]
$52.3$
$-10.5$
$61.4$
$-7.4$
$32.8$
$-9.2$
$32.6$
$-3.3$
MM2D3D [5]
$66.8$
$+4.0$
$70.2$
$+1.4$
$50.3$
$+8.3$
$46.1$
$+10.2$
Ours
$\mathbf{73.6}$
$\mathbf{+10.8}$
$\mathbf{70.5}$
$\mathbf{+1.7}$
$\mathbf{64.9}$
$\mathbf{+22.9}$
$\mathbf{52.1}$
$\mathbf{+16.2}$
Oracle
$76.0$
$69.2$
$78.4$
$71.9$
Table 2: Results under four domain shift scenarios with ${}^{\circ}$ point cloud, where not all the points are visible in the images. We report the 3D network performance in terms of mIoU. Note that the 3D backbone in these experiments is MinkUNet32 [9].
Method nuScenes normal-$\rightarrow$ Sem.KITTI Sem.KITTI normal-$\rightarrow$ nuScenes nuScenes normal-$\rightarrow$ Waymo Waymo normal-$\rightarrow$ nuScenes
Source only
$27.7$
$+0.0$
$28.1$
$+0.0$
$29.4$
$+0.0$
$21.8$
$+0.0$
PL [30]
$30.0$
$+2.3$
$29.0$
$+0.9$
$31.9$
$+2.5$
$22.3$
$+0.5$
CosMix [37]
$30.6$
$+2.9$
$29.7$
$+1.6$
$31.5$
$+2.1$
$30.0$
$+8.2$
MM2D3D [5]
$30.4$
$+2.7$
$31.9$
$+3.8$
$31.3$
$+1.9$
$33.5$
$+11.7$
$32.9$
$+5.2$
$33.7$
$+5.6$
$34.1$
$+4.7$
$37.5$
$+15.7$
Ours
$\mathbf{48.5}$
$\mathbf{+20.8}$
$\mathbf{42.9}$
$\mathbf{+14.8}$
$\mathbf{44.9}$
$\mathbf{+15.5}$
$\mathbf{48.2}$
$\mathbf{+26.4}$
Oracle
$70.3$
$78.3$
$79.9$
$78.3$

### v4

@@ Abstract
Unsupervised domain adaptation (UDA) in 3D segmentation tasks presents a formidable challenge, primarily stemming from the sparse and unordered nature of point clouds. Especially for LiDAR point clouds, the domain discrepancy becomes obvious across varying capture scenes, fluctuating weather conditions, and the diverse array of LiDAR devices in use. Inspired by the remarkable generalization capabilities exhibited by the vision foundation model, SAM, in the realm of image segmentation, our approach leverages the wealth of general knowledge embedded within SAM to unify feature representations across diverse 3D domains and further solves the 3D domain adaptation problem. Specifically, we harness the corresponding images associated with point clouds to facilitate knowledge transfer and propose an innovative hybrid feature augmentation methodology, which enhances the alignment between the 3D feature space and SAM’s feature space, operating at both the scene and instance levels. Our method is evaluated on many widely-recognized datasets and achieves state-of-the-art performance.
Keywords:
Unsupervised Domain Adaptation 3D Segmentation Feature Alignment Vision Foundation Model
@@ 1 Introduction
Figure 1: (a) Comparison of 3D UDA paradigms. Different from aligning two point feature domains directly, our method makes both the source domain and target domain align with the SAM feature space. (b) Visualization of the feature distance across different datasets, where smaller values indicate a more similar distribution. It is obvious that after mapping to SAM feature space, point feature distributions from disparate domains become much more aligned.
@@
Considering that RGB cameras yield dense, color-rich, and structured data, and more importantly, they represent minor discrepancies across various devices, certain 3D UDA methods [19, 5, 4] utilize the synergy of LiDAR and camera capabilities to achieve more comprehensive and precise perception, and further enhance adaptation capabilities for 3D segmentation tasks. However, these methods usually train 2D and 3D networks simultaneously, demanding substantial online computing resources. Vision foundation models (VFMs), such as the Segment Anything Model (SAM) [23], have garnered significant attention due to their remarkable performance in addressing open-world vision tasks. Such models are trained on massive image data with tremendous parameters. Compared with a common model trained on limited data, VFMs have more general knowledge and much stronger generalization capability. Many works such as  [8, 7] have emerged recently to transfer the general 2D vision knowledge of VFMs to 3D and have achieved promising performance.
@@
Based on SAM, focusing on image segmentation, we propose a novel paradigm for 3D UDA segmentation. As shown in Fig. 1(a), different from previous UDA approaches that strive to align the target domain to the source domain so that the model trained on labeled source data can also work on target data without annotation, our method makes both the source domain and target domain align with the SAM feature space. SAM feature space contains more general knowledge, which provides a friendly space to unify the feature representation from different domains. We utilize RGB images to assist point clouds in our framework. However, unlike the methods mentioned above only using images to provide auxiliary information, we take images as a bridge to align diverse 3D feature spaces to the SAM feature space, so we do not need to train extra 2D networks and we can process the image offline for less computing resources. Moreover, considering that the 3D feature space created by the source-domain data and the target-domain data is still much smaller than the SAM feature space, we propose a hybrid feature augmentation method at both scene and instance levels to generate more 3D data with diverse feature patterns in a broader data domain, which can further benefit the 3D-to-SAM feature alignment. In particular, we make full use of the masks generated by SAM to mix instance-level point clouds with the other domains. This technique can maintain the geometric completeness of instances, which is beneficial for semantic recognition.
@@
To verify that our idea of SAM-guided UDA is reliable, we randomly choose data from the source and target datasets and calculate the differences of feature distributions from source and target domain by KL divergence. The qualitative results are shown in Fig. 1(b), where the feature distribution differences after mapping to SAM feature space truly become much smaller. Moreover, to verify the effectiveness of our method, we compare it with current SOTA works on extensive 3D UDA segmentation settings and our method outperforms others by a large margin, improving about $14\%$ mIoU for VirtualKITTI-to-SematicKITTI, about $15\%$ mIoU for Waymo-to-nuScenes, and about $20\%$ mIoU for nuScenes-to-SemanticKITTI domain adaptation. Surprisingly, our unsupervised method achieves comparable performance with the supervised method for city-changing and light-changing settings on the nuScenes dataset. We also test our method on more challenging tasks, such as panoptic segmentation and domain generalization, showing that our method is robust and has good generalization capability.
@@ 2.2 Domain Adaptation for 3D Segmentation
Unsupervised Domain Adaptation (UDA) aims at transferring knowledge learned from a source annotated domain to a target unlabelled domain, and there are already several UDA methods proposed for 2D segmentation [6, 47, 22, 50]. In recent years, domain adaptation techniques have gained increasing traction in the context of 3D segmentation tasks. [46] leverage a "Complete and Label" strategy to enhance semantic segmentation of LiDAR point clouds by recovering underlying surfaces and facilitating the transfer of semantic labels across varying LiDAR sensor domains. CosMix [38] introduces a sample mixing approach for UDA in 3D segmentation, which stands as the pioneering UDA approach utilizing sample mixing to alleviate domain shift. It generates two new intermediate domains of composite point clouds through a novel mixing strategy applied at the input level, mitigating domain discrepancies. However, due to the sparsity and irregularity of the point cloud, the disparity across different point cloud data domains is larger compared to that across 2D image domains, which makes it difficult to mitigate the variation across domains.
@@ 2.3 Vision Foundation Models
The rise of foundation models [12, 21, 40] has garnered significant attention which are trained on extensive datasets, consequently demonstrating exceptional performance. Foundation models [49, 42] have seen significant advancements in the realm of 2D vision, and several research studies extend these foundation models to comprehend 3D information. Representative works CLIP [36] leverage contrastive learning techniques to train both text and image encoders. CLIP2Scene [8] extends the capabilities of CLIP by incorporating a 2D-3D calibration matrix to facilitate a deeper comprehension of 3D scenes and OpenScene[33] focuses on zero-shot learning for 3D scenes through aligning point features in CLIP feature space to enable open vocabulary queries for 3D points. The Meta Research team recently launched the ‘Segment Anything Model’ [23], trained on an extensive dataset of over 1 billion masks from 11 million images. Utilizing efficient prompting, SAM can generate high-quality masks for image instance segmentation. The integration of flexible prompting and ambiguity awareness enables SAM with robust generalization capabilities for various downstream segmentation challenges. Many methods [7, 8, 28] take it as an off-the-shelf tool and distillate the knowledge to solve 3D problems by 2D-3D feature alignment. In our work of tackling the UDA of 3D segmentation, we utilize SAM to provide 2D prior knowledge for 3D feature alignment in a wider data domain.
@@ 3.1 Problem Statement
We explore UDA for 3D segmentation, in which we have the source domain, denoted as $D_{S}=\{P_{S},I_{S},Y_{S}\}$ with paired input, namely point cloud $P_{S}$ and image $I_{S}$, as well as annotated labels $Y_{S}$ for each point, and the target domain denoted as $D_{T}=\{P_{T},I_{T}\}$ without any annotation. Using these data, we train a 3D segmentation model that can generalize well to the target domain. 3D data from various domains have obvious differences in distribution and patterns, leading to over-fitting problems when models trained in one domain try to analyze data from another. The main solution is to align different features despite domain differences to achieve the generalization capability of the model. Different from previous works, our novel paradigm is to map data from distinct domains into a unified feature space, ensuring the model performs consistently across domains.
Figure 2: Pipeline of our method. The point cloud is fed into the point encoder for point embeddings at the top, and the corresponding images are passed through the SAM encoder for image embeddings at the bottom, from which we obtain SAM-guided point embedding with the 2D-3D projection. Alignment loss $L_{align}$ is calculated based on the SAM-guided features and original features. Furthermore, augmented inputs provide diverse feature patterns boosting the 3D-to-SAM feature alignment.
@@ 3.2 Framework Overview
The vision foundation model, SAM, is trained by massive image data, which contains general vision knowledge and provides a friendly feature space to unify diverse feature representations. Taking 2D images as the bridge, the 3D feature space of different domains can be indirectly unified by bringing them closer to the SAM feature space based on 2D-to-3D knowledge distillation. Based on this, we design a novel SAM-guided UDA method for 3D segmentation, as Fig. 2 shows. Specifically, given a point cloud input $P$, the point encoder $M$ generates a point embedding $F_{point}\in\mathbb{R}^{n\times d}$ in the $d$-dimensional latent feature space. Concurrently, the corresponding image input $I$ is passed through the SAM encoder for a c-channels image embedding $F_{image}\in\mathbb{R}^{h\times w\times c}$. Utilizing the correspondence between the point cloud and image, we acquire SAM-guided point embedding $\hat{F}_{point}\in\mathbb{R}^{n\times d}$ to compute the alignment loss $L_{align}$ with the original point embedding $F_{point}$, serving the purpose of using SAM as a bridge to integrate the features of diverse data domains into a unified feature space. Notably, during training, the input for feature alignment consists of data from both source and target domains. We named this process as SAM-guided Feature Alignment. At the same time, as for labeled data $Y$, segmentation loss $L_{seg}$ is also calculated as semantic supervision. During model training, only the point cloud branch of the whole pipeline is trained, and the gradient is not calculated in the image branch, which makes our method more lightweight. Furthermore, a Scene-Instance Hybrid Feature Augmentation is designed, which consists of scene-level and instance-level mix-up strategy. These mix-up strategies boost the variance of training data and generalize the network capability under the convex combination of the source domain and target domain data. Notably, the instance-level feature augmentation could maintain the local geometric relationship between two domains and make the subsequent alignment efficient.
@@ 4.1 Dataset Setup
We first follow the benchmark introduced in xMUDA [20] to evaluate our method, comprehending four domain shift scenarios, including (1) USA-to-Singapore, (2) Day-to-Night, (3) VirtualKITTI-to-SemanticKITTI and (4) A2D2-to-Semantic-KITTI. The first two leverage nuScenes [3] as their dataset, consisting of 1000 driving scenes in total with 40k annotated point-wise frames. Specifically, the former differs in the layout and infrastructure while the latter exhibits severe illumination changes between the source and the target domain. The third is more challenging since it is the adaptation from synthetic to real data, implemented by adapting from VirtualKITTI [14] to SemanticKITTI [2] while the fourth involves A2D2 [15] and SemanticKITTI as different data domains, where the domain discrepancy lies in the distinct density and arrangement of 3D point clouds captured by different devices since the A2D2 is captured by 16-beam LiDAR and the SemanticKITTI uses 64-beam LiDAR. For the above settings, noted that only 3D points visible from the camera are used for training and testing, specifically, only one image and corresponding points for each sample are used for training.
Since we only use the image combined with SAM as offline assistance for the training of a 3D segmentation network instead of training a new 2D segmentation network, we focus on comparing the performance of the 3D segmentation network and enable model training with the whole point cloud sample because of less computational cost, even if some part of it is not visible in the images. For the part of the point cloud that cannot be covered by the image, alignment loss is not calculated, only segmentation loss is calculated. Thus, we also compare our method with others trained with the whole 360 view of the point cloud, in which three datasets are involved including nuScenes, SemanticKITTI, and Waymo [29]. In these settings, we use 6 images in nuScenes covering 360 view, 1 image in SemanticKITTI covering 120 view, and 5 images in Waymo covering 252 view. More information is introduced in the Appendix. For metric, We compute the Intersection over the Union per class and report the mean Intersection over the Union (mIoU).
@@ 4.2 Implementation Details
We make source and target labels compatible across these experiments. For all benchmarks in prior multi-modal UDA methods, we strictly follow class mapping like xMUDA for a fair comparison, while we map the labels of the dataset in other experiments into 10 segmentation classes in common. Our method is implemented by using the public PyTorch [31] repository MMDetection3D [11] and all the models are trained on a single 24GB GeForce RTX 3090 GPU. To compare fairly, we use SparseConvNet [16] with U-Net architecture as the 3D backbone network when following the benchmark introduced in xMUDA to evaluate our method and use MinkUNet32 [9] as the 3D backbone network when following the setup of taking the whole 360 point cloud as input, which is also the backbone of the state-of-the-art uni-modal method CosMix. For the image branch, the ViT-h variant SAM model is utilized to generate image embedding for SAM-guided feature alignment and instance masks for hybrid feature augmentation in an offline manner. We keep the proportion of mixed data and normal data from the source and target domain the same during model training. Before the data is fed into the 3D network, data augmentation such as vertical axis flipping, random scaling, and random 3D rotations are widely used like all the compared methods. For the model training strategies, we choose a batch size of 8 for both source data and target data, then mix the data batch for training at each iteration. Besides, we adopt AdamW as the model optimizer and One Cycle Policy as the learning-rate scheduler.
@@
Method USA $\rightarrow$ Singapore Day $\rightarrow$ Night v.KITTI $\rightarrow$ Sem.KITTI A2D2 $\rightarrow$ Sem.KITTI
Source only $62.8$ $+0.0$ $68.8$ $+0.0$ $42.0$ $+0.0$ $35.9$ $+0.0$
xMUDA [20] $63.2$ $+0.4$ $69.2$ $+0.4$ $46.7$ $+4.7$ $46.0$ $+10.1$
DsCML [32] $52.3$ $-10.5$ $61.4$ $-7.4$ $32.8$ $-9.2$ $32.6$ $-3.3$
MM2D3D [5] $66.8$ $+4.0$ $70.2$ $+1.4$ $50.3$ $+8.3$ $46.1$ $+10.2$
Ours $\mathbf{73.6}$ $\mathbf{+10.8}$ $\mathbf{70.5}$ $\mathbf{+1.7}$ $\mathbf{64.9}$ $\mathbf{+22.9}$ $\mathbf{52.1}$ $\mathbf{+16.2}$
Oracle $76.0$- $69.2$- $78.4$- $71.9$-
Table 2: Results under four domain shift scenarios with 360 point cloud, where not all the points are visible in the images. We report the 3D network performance in terms of mIoU. Note that the 3D backbone in these experiments is MinkUNet32 [9].
Method nuScenes $\rightarrow$ Sem.KITTI Sem.KITTI $\rightarrow$ nuScenes nuScenes $\rightarrow$ Waymo Waymo $\rightarrow$ nuScenes
Source only $27.7$ $+0.0$ $28.1$ $+0.0$ $29.4$ $+0.0$ $21.8$ $+0.0$
PL [30] $30.0$ $+2.3$ $29.0$ $+0.9$ $31.9$ $+2.5$ $22.3$ $+0.5$
CosMix [37] $30.6$ $+2.9$ $29.7$ $+1.6$ $31.5$ $+2.1$ $30.0$ $+8.2$
MM2D3D [5] $30.4$ $+2.7$ $31.9$ $+3.8$ $31.3$ $+1.9$ $33.5$ $+11.7$
MM2D3D $32.9$ $+5.2$ $33.7$ $+5.6$ $34.1$ $+4.7$ $37.5$ $+15.7$
Ours $\mathbf{48.5}$ $\mathbf{+20.8}$ $\mathbf{42.9}$ $\mathbf{+14.8}$ $\mathbf{44.9}$ $\mathbf{+15.5}$ $\mathbf{48.2}$ $\mathbf{+26.4}$
Oracle $70.3$- $78.3$- $79.9$- $78.3$-